OpenStudy (cutiecomittee123):

What functions are equivelent to y=4cosx-2 y=4cos(-x)-2 y=4sin(x+pi/2)-2 y=4sin(x-pi/2)-2 y=-4cosx+2

OpenStudy (jango_in_dtown):

cos(-x)=cos x . now try

imqwerty (imqwerty):

these are some identities try using them to get the answer \[\cos(-x)=\cos(x)\]\[\sin \left( \frac{ \pi }{ 2 }+x \right)=\sin(x)\]\[\sin \left( \frac{ \pi }{ 2 }-x \right)=\cos(x)\]

OpenStudy (cutiecomittee123):

so a and d work in that case?

imqwerty (imqwerty):

d won't work

OpenStudy (cutiecomittee123):

how so?

OpenStudy (cutiecomittee123):

is it because its not the x that is being negative, its the whole equation?

OpenStudy (cutiecomittee123):

and is that the only one that is equal?

imqwerty (imqwerty):

so if d is the answer -4cos(x)+2=4cos(x)-2 bring both on right hand side 0=8cos(x)-4 is this correct no its not so d won't wrk what do u think about c

OpenStudy (cutiecomittee123):

oh i see, so if you just set them equal to eachother and solve then you will get your answers!!!!!

imqwerty (imqwerty):

yea u can check the options like that

imqwerty (imqwerty):

if the option u choose satisfies the equation then its correct :D

OpenStudy (cutiecomittee123):

cool thanks

jimthompson5910 (jim_thompson5910):

You can also use a graphing calculator to check. https://www.desmos.com/calculator/lwlzconupi Notice how the original graph of `y=4cos(x)-2` (in red) does not match with choice D `y=-4cos(x)+2` (in blue). If they did match, then one curve would be right on top of the other. You can turn on/off the graphs to see if one was on top of the other.

jimthompson5910 (jim_thompson5910):

this shows how the original graph matches up with choice A https://www.desmos.com/calculator/gwwpecq6te click the circle icons to turn the graph on/off and you'll see one graph overlapping perfectly on the other

OpenStudy (cutiecomittee123):

well b and c cannot match because they are sin and the origional is cosine, right?

OpenStudy (cutiecomittee123):

actually i just tested it and b actually matches up

jimthompson5910 (jim_thompson5910):

btw @imqwerty it should be sin(pi/2 + x) = cos(x)

OpenStudy (cutiecomittee123):

https://www.desmos.com/calculator/lwlzconupi

jimthompson5910 (jim_thompson5910):

yeah B should match with the original

OpenStudy (cutiecomittee123):

so b and d are both equal

imqwerty (imqwerty):

>.< oh srry silly mistake

jimthompson5910 (jim_thompson5910):

`so b and d are both equal` incorrect

OpenStudy (cutiecomittee123):

yeah but b and d match up

OpenStudy (cutiecomittee123):

so then just d? was I right about the fact that they cannot be the same because b and c are sine and the origional is cosine?

jimthompson5910 (jim_thompson5910):

since sin(pi/2 + x) = cos(x), this means y=4sin(x+pi/2)-2 turns into y=4cos(x)-2

OpenStudy (cutiecomittee123):

okay i kind of get that so does that mean that c works

jimthompson5910 (jim_thompson5910):

one moment

OpenStudy (cutiecomittee123):

i graphed it and they matched up

jimthompson5910 (jim_thompson5910):

ok I graphed the original equation (in box 1) and the four answer choices (boxes 2 through 5). Only box 1 is turned on. The other graphs are turned off https://www.desmos.com/calculator/koejh8yigm I recommend going through each and turn them on one at a time. example: turn on box 1 and box 3 to compare the original with choice B

jimthompson5910 (jim_thompson5910):

hopefully you'll be able to see that A matches B matches C does not match D does not match

OpenStudy (cutiecomittee123):

yep that what i see too. I tested it out

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (cutiecomittee123):

sweet thanks